OpenStudy (anonymous):

Felipe is going for a run. He runs for 6.4 miles at a speed of 4 miles per hour. For how many hours does he run? is the answer 2 hours?

OpenStudy (anonymous):

well technically its 1.6 but if you round up then yes it is 2 hrs

OpenStudy (anonymous):

Solve the following proportion for y. y/5 =11/3 round your answer to the nearest tenth

OpenStudy (anonymous):

u wanna give me a medal first? :)

OpenStudy (anonymous):

sorry wasnt paying attention

OpenStudy (anonymous):

k thanks! lets move on shall we? ok lemme draw the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ y }{ 5 }=\frac{ 11 }{ 3 }\]

OpenStudy (anonymous):

so cross multiply do you know how to do that?

OpenStudy (anonymous):

3y/55 =55/3y?

OpenStudy (anonymous):

no not quite here let me draw it out

OpenStudy (anonymous):